OpenStudy (anonymous):

how do you take the function y=x^3+5|x| and break it into two separate functions to avoid the corner at (0,0)?

OpenStudy (anonymous):

If x<0,\[\left| x \right|=-x\] if x>0, \[\left| x \right|=x\] hence functions are y=x^3-5x and y=x^3+5x